# Diffusion Models, Image Super-Resolution And Everything: A Survey

### v2

@@ V-D2 Classifier-Free Guidance
V-E Zero-Shot
Zero-shot image SR aims to develop methods that do not depend on prior image examples or training [119, 16]. Typically, these methods harness the inherent redundancy within a single image for improvement. They often leverage pre-trained DMs for generation, incorporating LR images as conditions during the sampling process, in contrast to other conditioning methods discussed earlier. Additionally, they differ from guidance-based methods, where conditioning information is used to weight the training of a DM from scratch. A recent study by Li et al. [113] categorizes diffusion-based methods into projection-based, decomposition-based, and posterior estimation, which are introduced in this section. The discussed methods are compared in Table IV.
V-E1 Projection-Based
@@
Figure 6: Overview of DDNM [117]. It utilizes the range-null space decomposition for constructing a general solution for multiple tasks, such as image SR, colorization, inpainting, and deblurring.
V-E2 Decomposition-Based
@@
[…]
@@
The last step involves defining $\mathbf{A}$ and $\mathbf{A^{\dagger}}$, the construction of which is contingent on the restoration task at hand. For instance, in SR tasks involving scaling by a factor of $n$, $\mathbf{A}$ can be defined as a $1\times n^{2}$ matrix, representative of an average-pooling operator. The average-pooling operator, denoted as $\begin{bmatrix}\frac{1}{n^{2}}&...&\frac{1}{n^{2}}\end{bmatrix}$, functions to average each patch into a singular value. Similarly, we can construct its pseudo-inverse as $\mathbf{A}^{\dagger}\in\mathbb{R}^{n^{2}\times 1}=\begin{bmatrix}1&...&1\end{bmatrix}^{\top}$. The original work provides further examples of tasks (such as colorization, inpainting, and restoration), illustrating how these methods are applied. In addition, it describes how compound operations consisting of numerous sub-operations function in these contexts. In their research, the authors also introduced ${}^{+}$ to support the restoration of noisy images. They utilized a technique analogous to the ”back and forward” strategy implemented in RePaint [120]. This approach was leveraged to enhance the quality further.
Given this approach’s novelty, only a handful of subsequent studies extend and build upon it, such as the work presented in CDPMSR [92]. This research direction promises exciting possibilities for future exploration, although it calls for further investigation. For example, it should be noted that the DDNM approach introduces additional computational expenses compared to the task-specific training carried out using DDPMs. Moreover, the degradation operator $\mathbf{A}$ is set manually, which can be challenging for certain tasks. Another potential drawback is the assumption that $\mathbf{A}$ functions as a linear degradation operator, which may not always hold true and thus could limit the model’s effectiveness in certain scenarios.
V-E3 Posterior Estimation
@@
[…]
V-F Corruption Space
Karras et al. [128] identified three pillars of DMs: the noise schedule, the network parameterization, and the sampling algorithm. Recently, many authors argued to consider also different types of corruption used during forward diffusion like Soft Score Matching [129]. It directly incorporates the filtering process within the SGM, training the model to predict a clean image. Upon corruption, this predicted image aligns with the diffused observation.
Cold Diffusion [130] presents another ingenious way of modifying the corruption space for DDPMs, which shows that the generative capability is not strongly dependent on the choice of image degradation. It reveals new experimental types of diffusion besides Gaussian noise can be effectively used, like animorphosis (i.e., human faces iteratively degrading to animal faces).
[…]
Another avenue for alternative corruption space is presented by Inversion by Direct Iteration (InDI) [132]. InDI delineates a direct mapping strategy, efficiently bridging the gap between the two quality spaces without the iterative refinement typically required by conventional diffusion processes. The intrinsic flexibility and the direct mapping capability of InDI propose intriguing possibilities for enhancing image quality, suggesting a potent avenue for research exploration. The potential integration of InDI’s principles with those of conditional diffusion models could offer substantial advancements in the field of image SR. It is anticipated that a detailed examination and discussion of InDI within the broader scope of diffusion-based image enhancement could yield valuable insights and contribute significantly to the ongoing development of generative models in image processing.
V-G Color Shifting
V-H Architecture Designs for Denoising
The design of the denoising model in DMs offers a range of options. The majority of DMs adopt the use of U-Net, as noted in most literature [110]. SR3 [15], for instance, employs residual blocks from BigGAN [110] and re-scales skip connections by a factor of $\frac{1}{\sqrt{2}}$. SRDiff takes a similar approach [76], although it opts for vanilla residual blocks without the re-scaling of skip connections and uses a LR encoder to incorporate the information of the LR image during the backward diffusion. Whang et al. [88] exploit an initial predictor to combine the strengths of deterministic image SR models and DMs, as shown in Figure 9. It has the advantage that the DM only needs to learn the residuals that the deterministic image SR model (initial predictor) fails to predict, hence simplifying the learning target. Additionally, the removal of self-attention, positional encodings and group normalization from the SR3 U-Net enables their model to support arbitrary resolutions. An initial predictor is also employed in the wavelet-based approach DiWa [84]. Moreover, wavelet SR models, such as DWSR [83] – a simple sequence of convolution layers of depth 10 – are utilized for denoising prediction in the wavelet domain. In WaveDM [85], a deterministic U-Net predictor is used for the high-frequency band, while diffusion is applied in the low-frequency band.
Figure 9: Illustration of an initial predictor like in Whang et al. [88]. It divides the computation into a deterministic path (initial predictor) and a stochastic part (denoiser network), which predicts the missing residual information.
VI Domain-Specific Applications
@@
SR3 [15] produces photo-realistic and perceptually state-of-the-art images on faces and natural images but may not be suitable for other tasks like remote sensing. Some models are more suited to certain tasks as they tackle issues specific to the domain. This section highlights the applications of DMs to domain-specific SR tasks: Medical imaging, special cases of face SR (Blind Face Restoration and Atmospheric Turbulences), and remote sensing.
VI-A Medical Imaging
@@
Magnetic Resonance Imaging (MRI) scans are widely used to aid patient diagnosis but can often be of low quality and corrupted with noise. Chung et al. [137] propose a combined denoising and SR network referred to as R2222D2+limit-from22+2 + (Regularized Reverse Diffusion Denoiser $+$ SR). They perform denoising of the MRI scans, followed by an SR module. Inspired by CCDF (i.e., a zero-shot method) from Chung et al. [123], they start their backward diffusion from an initial noisy image instead of pure Gaussian noise. The reverse SDE is solved using a non-parametric, eigenvalue-based method. In addition, they restrict the stochasticity of the DMs through low-frequency regularization. Particularly, they maintain low-frequency information while correcting the high-frequency ones to produce sharp and super-resolved MRI scans. Mao et al. [138] addresses the lack of diffusion-based multi-contrast MRI SR methods. They propose a Disentangled Conditional Diffusion model (DisC-Diff) to leverage a multi-conditional fusion strategy based on representation disentanglement, enabling high-quality HR image sampling. Specifically, they employ a disentangled U-Net with multiple encoders to extract latent representations and use a novel joint disentanglement and Charbonnier loss function to learn representations across MRI contrasts. They also implement curriculum learning and improve their MRI model for varying anatomical complexity by gradually increasing the difficulty of training images.
VI-B Blind Face Restoration
@@
A solution to BFR was proposed by Zongsheng et al. with DifFace [139] that leverages the rich generative priors of pre-trained DMs with parameters $\theta$, which were trained to approximate $p_{\theta}(\mathbf{z}_{t}|\mathbf{z}_{t-1})$. In contrast to existing methods that learn direct mappings from $\mathbf{x}$ to $\mathbf{y}$ under several constraints [140, 141], DifFace circumvents this by generating a diffused version $\mathbf{z}_{N}$ of the desired HR image $\mathbf{y}$ with $N<T$. They predict the starting point, the posterior $q(\mathbf{z}_{N}|\mathbf{x})$ via a transition distribution $p(\mathbf{z}_{N}|\mathbf{x})$. The transition distribution is formulated like the regular diffusion process, a Gaussian distribution, but uses an initial predictor $\varphi(\mathbf{x})$ to generate the mean, named diffused estimator. As their model borrows the reverse Markov chain from a pre-trained DM, DifFace requires no full retraining for new and unknown degradations, unlike SR3.
@@
VI-C Atmospheric Turbulence in Face SR
@@
VI-D Remote Sensing
Remote Sensing Super-Resolution (RSSR) addresses the HR reconstruction from one or more LR images to aid object detection and semantic segmentation tasks for satellite imagery. RSSR is limited by the absence of small targets with complex granularity in the HR images [146]. To produce finer details and texture, Liu et al. [147] present DMs with a Detail Complement mechanism (DMDC). They train their model similar to SR3 [15] and perform a detailed supplement task. To generate high-frequency information, they randomly mask several parts of the images to mimic dense objects. The SR images recover the occluded patches as the model learns small-grained information. Additionally, they introduce a novel pixel constraint loss to limit the diversity of DMDC and improve overall accuracy. Ali et al. [148] design a new architecture for RS images that integrates Vision Transformers (ViT) with DMs as a Two-stage approach for Enhancement and Super-Resolution (TESR). In the first stage (SR stage), the SwinIR [108] model is used for RSSR. In the second stage (enhancement stage), the noisy images are enhanced by employing DMs to reconstruct the finer details. Xu et al. [149] propose a blind SR framework based on Dual conditioning DDPMs for SR (DDSR). A kernel predictor conditioned on LR image encodings estimates the degradation kernel in the first stage. This is followed by an SR module consisting of a conditional DDPM in a U-Net with the predicted kernel and the LR encodings as guidance. An RRDB encoder extracts the encodings from LR images.
VII Discussion and Future Work
@@ VII-A Color Shifting
Often, the most practical advancements come from a solid theoretical understanding. As discussed in subsection V-G, due to the substantial computational demands, DMs may occasionally exhibit color shifts when constrained by hardware limitations that demand smaller batch sizes or shorter training periods [133]. While well-defined diffusion methods [121] or color normalization [100] might mitigate this problem, a theoretical understanding of why it is emerging is necessary.
VII-B Computational Costs
@@
In a study conducted by Ganguli et al., it was observed that the computing power needed for large-scale AI experiments has surged by over 300,000 times in the last decade [150]. Regrettably, this increase in resource intensity has been accompanied by a sharp decline in the share of these results originating from academic circles. DMs are not immune to this issue; their computational demands add to the expanding gap between industry and academia. Therefore, there is a pressing need to reduce computational costs and memory footprints for practical applicability and research. One strategy to alleviate computational demands is to examine smaller spatial-sized domains, as discussed in subsection V-B. Examples of such approaches include LDMs [13, 9] and wavelet-based models [84, 86]. However, the capability of LDMs to reconstruct data with high precision and fine-grained accuracy, as required in image SR, remains to be questioned. Therefore, further advancements in these methods are critically needed. On the other hand, wavelet-based models do not present a bottleneck regarding information preservation. This advantage suggests that they should be the subject of more intensive exploration.
VII-C Efficient Sampling
[…]
VII-D Corruption Spaces
[…]
VII-E Comparability
@@
The comparison of different DMs in image SR is challenging due to the use of varied datasets across studies. These datasets differ in aspects like resolution, content diversity, color distribution, and noise levels, affecting a model’s performance. DM might excel with one dataset but not another, making it difficult to gauge its overall effectiveness. Thus, the performance of a DM on a specific dataset does not necessarily reflect its general effectiveness. To address this comparability issue, it is crucial to establish a standard benchmark consisting of diverse, representative datasets and a standardized set of evaluation metrics, not only for DM-based methods. This benchmark should cover various degradation types and levels, allowing a fair and comprehensive comparison of DMs in image SR. A well-structured benchmark can facilitate the identification of models that perform well across various tasks and conditions, thus accelerating the advancement of the field. Also, assessing the quality of SR images produced by generative models remains a challenging and unresolved issue. While DMs underperform on standard metrics (PSNR and SSIM), the images generated are more photo-realistic, unlike the results obtained from classical image SR works [16]. Consequently, generative models like DMs tend to show lower scores using these metrics compared to pixel-based methods, despite often being rated more favorably by human evaluators [15]. LPIPS [155] performs better reflecting this perception, but the domain of image SR has to adapt to more diverse metrics, such as predictors that reflect human ratings directly [156, 157]. For instance, datasets with subjective ratings, like TID2013 [158], and neural networks, such as DeepQA [159] or NIMA [160], can be employed to predict human-like scoring of images and should be further explored in the context of diffusion generated SR images.
VII-F Image Manipulation
@@
The authors of SRDiff [76], proposed two potential extensions of SR DMs: content fusion and latent space interpolation. Content fusion involves the combination of content from two source images. For instance, they replace the eyes in one source image with the face from another image before conducting diffusion in the image space like CutMix [161]. The backward diffusion successfully creates a smooth transition between both images. In the latent space interpolation model, the latent space of two SR predictions is linearly interpolated to generate a new image. While these extensions have yielded remarkable results, unlike other generative models such as Variational Autoencoders (VAEs) or Generative Adversarial Networks (GANs), DMs have been found to offer less proficient latent representations [162]. Therefore, recent and ongoing research into the manipulation of latent representations in diffusion models is both in its early stages and greatly needed [163, 164, 165]. This can be particularly useful in multi-image SR for generating high-resolution images that blend characteristics from multiple sources, potentially improving the quality and diversity of the output (e.g., satellite imagery for SR predictions with flexible daylights).
VII-G Cascaded Image Generation

### v3

@@ V-G Architecture Designs for Denoising
VI Diffusion-based Zero-Shot SR
Zero-shot image SR aims to develop methods that do not depend on prior image examples or training [153, 16]. Typically, these methods harness the inherent redundancy within a single image for improvement. They often leverage pre-trained DMs for generation, incorporating LR images as conditions during the sampling process, in contrast to other conditioning methods discussed earlier [154]. Additionally, they differ from guidance-based methods, where conditioning information is used to weight the training of a DM from scratch. A recent study by Li et al. [17] categorizes diffusion-based methods into projection-based, decomposition-based, and posterior estimation, which are introduced in this section. The discussed methods are compared in Table IV.
VI-A Projection-Based
@@
VI-B Decomposition-Based
@@
$\textit{Consistency}:\quad\mathbf{A}\mathbf{\hat{y}}\equiv\mathbf{x},\quad\quad\textit{Realness}:\quad\mathbf{\hat{y}}\sim p(\mathbf{y}),$ (39)
@@
The last step involves defining $\mathbf{A}$ and $\mathbf{A^{\dagger}}$, the construction of which is contingent on the restoration task at hand. For instance, in SR tasks involving scaling by a factor of $n$, $\mathbf{A}$ can be defined as a $1\times n^{2}$ matrix, representative of an average-pooling operator. The average-pooling operator, denoted as $\begin{bmatrix}\frac{1}{n^{2}}&...&\frac{1}{n^{2}}\end{bmatrix}$, functions to average each patch into a singular value. Similarly, we can construct its pseudo-inverse as $\mathbf{A}^{\dagger}\in\mathbb{R}^{n^{2}\times 1}=\begin{bmatrix}1&...&1\end{bmatrix}^{\top}$. The original work provides further examples of tasks (such as colorization, inpainting, and restoration), illustrating how these methods are applied. In addition, it describes how compound operations consisting of numerous sub-operations function in these contexts. In their research, the authors also introduced DDNM+ to support the restoration of noisy images. They utilized a technique analogous to the ”back and forward” strategy implemented in RePaint [155]. This approach was leveraged to enhance the quality further.
Given this approach’s novelty, only a handful of subsequent studies extend and build upon it, such as the work presented in CDPMSR [121]. This research direction promises exciting possibilities, although it calls for further investigation. For example, it should be noted that the DDNM approach introduces additional computational expenses compared to the task-specific training carried out using DDPMs. Moreover, the degradation operator $\mathbf{A}$ is set manually, which can be challenging for certain tasks. Another potential drawback is the assumption that $\mathbf{A}$ functions as a linear degradation operator, which may not always hold true and thus could limit the model’s effectiveness in certain scenarios.
Figure 8: Overview of DDNM [151]. It utilizes the range-null space decomposition to construct a general solution for multiple tasks, such as image SR, colorization, inpainting, and deblurring.
VI-C Posterior Estimation
@@
VII Domain-Specific Applications
@@
SR3 [15] produces photo-realistic and perceptually state-of-the-art images on faces and natural images but may not be suitable for other tasks like remote sensing. Some models are more suited to certain tasks as they tackle issues specific to the domain [161]. This section highlights the applications of DMs to domain-specific SR tasks: Medical imaging, special cases of face SR (Blind Face Restoration and Atmospheric Turbulences), and remote sensing.
VII-A Medical Imaging
@@
Magnetic Resonance Imaging (MRI) scans are widely used to aid patient diagnosis but can often be of low quality and corrupted with noise. Chung et al. [162] propose a combined denoising and SR network referred to as R2222D2+limit-from22+2 + (Regularized Reverse Diffusion Denoiser $+$ SR). They perform denoising of the MRI scans, followed by an SR module. Inspired by CCDF (i.e., a zero-shot method) from Chung et al. [156], they start their backward diffusion from an initial noisy image instead of pure Gaussian noise. The reverse SDE is solved using a non-parametric, eigenvalue-based method. In addition, they restrict the stochasticity of the DMs through low-frequency regularization. Particularly, they maintain low-frequency information while correcting the high-frequency ones to produce sharp and super-resolved MRI scans. Mao et al. [163] addresses the lack of diffusion-based multi-contrast MRI SR methods. They propose a Disentangled Conditional Diffusion model (DisC-Diff) to leverage a multi-conditional fusion strategy based on representation disentanglement, enabling high-quality HR image sampling. Specifically, they employ a disentangled U-Net with multiple encoders to extract latent representations and use a novel joint disentanglement and Charbonnier loss function to learn representations across MRI contrasts. They also implement curriculum learning and improve their MRI model for varying anatomical complexity by gradually increasing the difficulty of training images. An improvement of DisC-Diff by combining the DM with a transformer was introduced by Li et al. with DiffMSR [164].
VII-B Blind Face Restoration
@@
A solution to BFR was proposed by Yue et al. with DifFace [165] that leverages the rich generative priors of pre-trained DMs with parameters $\theta$, which were trained to approximate $p_{\theta}(\mathbf{z}_{t}|\mathbf{z}_{t-1})$. In contrast to existing methods that learn direct mappings from $\mathbf{x}$ to $\mathbf{y}$ under several constraints [166, 167], DifFace circumvents this by generating a diffused version $\mathbf{z}_{N}$ of the desired HR image $\mathbf{y}$ with $N<T$. They predict the starting point, the posterior $q(\mathbf{z}_{N}|\mathbf{x})$ via a transition distribution $p(\mathbf{z}_{N}|\mathbf{x})$. The transition distribution is formulated like the regular diffusion process, a Gaussian distribution, but uses an initial predictor $\varphi(\mathbf{x})$ to generate the mean, named diffused estimator. As their model borrows the reverse Markov chain from a pre-trained DM, DifFace requires no full retraining for new and unknown degradations, unlike SR3.
@@
VII-C Atmospheric Turbulence in Face SR
@@
VII-D Remote Sensing
Remote Sensing Super-Resolution (RSSR) addresses the HR reconstruction from one or more LR images to aid object detection and semantic segmentation tasks for satellite imagery. RSSR is limited by the absence of small targets with complex granularity in the HR images [172]. To produce finer details and texture, Liu et al. [173] present DMs with a Detail Complement mechanism (DMDC). They train their model similar to SR3 [15] and perform a detailed supplement task. To generate high-frequency information, they randomly mask several parts of the images to mimic dense objects. The SR images recover the occluded patches as the model learns small-grained information. Additionally, they introduce a novel pixel constraint loss to limit the diversity of DMDC and improve overall accuracy. Ali et al. [174] design a new architecture for RS images that integrates Vision Transformers (ViT) with DMs as a Two-stage approach for Enhancement and Super-Resolution (TESR). In the first stage (SR stage), the SwinIR [49] model is used for RSSR. In the second stage (enhancement stage), the noisy images are enhanced by employing DMs to reconstruct the finer details. Xu et al. [175] propose a blind SR framework based on Dual conditioning DDPMs for SR (DDSR). A kernel predictor conditioned on LR image encodings estimates the degradation kernel in the first stage. This is followed by an SR module consisting of a conditional DDPM in a U-Net with the predicted kernel and the LR encodings as guidance. An RRDB encoder extracts the encodings from LR images. Recently, Khanna et al. introduced DiffusionSat [176], which uses a LDM for RSSR and incorporates additional remote sensing conditioning information (e.g., longitude, latitude, cloud cover, etc.).
VIII Discussion and Future Work
@@ VIII-A Color Shifting
Often, the most practical advancements come from a solid theoretical understanding. As discussed in subsection V-F, due to the substantial computational demands, DMs may occasionally exhibit color shifts when constrained by hardware limitations that demand smaller batch sizes or shorter training periods [142]. While well-defined diffusion methods [143] or color normalization [107] might mitigate this problem, a theoretical understanding of why it is emerging is necessary.
VIII-B Computational Costs
@@
In a study conducted by Ganguli et al., it was observed that the computing power needed for large-scale AI experiments has surged by over 300,000 times in the last decade [177]. Regrettably, this increase in resource intensity has been accompanied by a sharp decline in the share of these results originating from academic circles. DMs are not immune to this issue; their computational demands add to the expanding gap between industry and academia. Therefore, there is a pressing need to reduce computational costs and memory footprints for practical applicability and research. One strategy to alleviate computational demands is to examine smaller spatial-sized domains, as discussed in subsection V-C. Examples of such approaches include LDMs [13, 5] and wavelet-based models [113, 115]. However, the capability of LDMs to reconstruct data with high precision and fine-grained accuracy, as required in image SR, remains to be questioned. Therefore, further advancements in these methods are critically needed. On the other hand, wavelet-based models do not present a bottleneck regarding information preservation. This advantage suggests that they should be the subject of more intensive exploration.
VIII-C Efficient Sampling
A benefit of DMs is the possibility of decoupling training and inference schedules [178]. This allows for substantial enhancements in curtailing the time required for inference in practical applications, providing a significant efficiency edge in real-world scenarios. While reducing the number of steps taken during inference is relatively simple, a systematic method for determining inference schedules has yet to be developed [179]. As outlined in subsection IV-A, this research direction represents a promising avenue. We explored training-based sampling methods for SR with AddSR [88] and YONOS-SR [89] but also introduced efficient DMs that need fewer sampling steps, like ResShift [118] and DiffIR [145]. An alternative is given by methods that use different corruption spaces, as discussed in subsection V-E. Unlike sampling from pure Gaussian noise, notable works such as Luo et al. [108], I2SB [140], Come-Closer-Diffuse-Faster [156], or Cold Diffusion [139] define a process from the LR to the HR image directly. Additional techniques for decreasing computation time, such as knowledge distillation, alternative noise schedulers, or truncated diffusion, demand further investigation concerning image SR [180, 181, 84, 85].
VIII-D Corruption Spaces
New approaches for corruption spaces allow a more direct approach for upsampling images from LR to HR. The significance of exploring different corruption spaces lies in addressing the inherent limitations and assumptions embedded within current DM frameworks, e.g., diversity and blurriness added during the forward diffusion process. The adaptability and efficiency demonstrated by novel approaches like InDI or I2SB, especially in handling diverse and complex corruption patterns, spotlight the urgent need for future research.
VIII-E Comparability
@@
Comparing DMs in SR is complex because of the varied datasets used in different studies. They vary in resolution, content diversity, color distribution, and noise levels, all of which significantly influence model performance. A model may perform well with one dataset but poorly with another, complicating the assessment of its overall effectiveness. Establishing a standard benchmark with diverse, representative datasets and uniform evaluation metrics is essential for comparability. This approach would help identify models that consistently perform well across different conditions and tasks, thereby promoting faster progress in the field. Furthermore, evaluating the quality of SR images from generative models is still problematic. Although DMs often produce more photorealistic images, they typically score lower on standard metrics like PSNR and SSIM [16]. However, these models tend to receive more favorable assessments from human evaluators [15]. LPIPS [182] performs better reflecting this perception, but the domain of image SR has to adapt to more diverse metrics, such as predictors that reflect human ratings directly [183, 184]. For instance, datasets with subjective ratings, like TID2013 [61], and neural networks, such as DeepQA [62] or NIMA [63], can be employed to predict human-like scoring of images and should be further explored.
VIII-F Image Manipulation
@@
Image manipulation can be particularly useful in multi-image SR for generating HR images that blend characteristics from multiple sources, potentially improving the quality and diversity of the output (e.g., satellite imagery for SR predictions with flexible daylights). SRDiff [79] proposed two potential extensions: content fusion and latent space interpolation. Content fusion involves the combination of content from two source images. For instance, they replace the eyes in one source image with the face from another image before conducting diffusion in the image space like CutMix [185]. The backward diffusion successfully creates a smooth transition between both images. In the latent space interpolation model, the latent space of two SR predictions is linearly interpolated to generate a new image. While these extensions have yielded remarkable results, unlike other generative models such as VAEs or GANs, DMs have been found to offer less proficient latent representations [186]. Therefore, recent and ongoing research into the manipulation of latent representations in DMs is both in its early stages and greatly needed [187, 188, 189].
VIII-G Cascaded Image Generation